OpenStudy (anonymous):

What is the solution of the system? use elimination. 3x - 4y = -24 x+y = -1

OpenStudy (anonymous):

hii

OpenStudy (here_to_help15):

what is the solution of the system? use elimination. 3x-4y=-24. x+y=-1? --------------------------------------... 3x-4y=-24 x = -y-1 substitute 3(-y-1) -4y = -24 expand and combine terms -3y -3 -4y = -24 add 3 to both sides -7y = -21 y = 3 substituting this into x = -y-1, we have x = -3-1, or x = -4 solution is (-4, 3).